OpenStudy (anonymous):

how would you solve for \[6^{2x} - 2(6^x)-15=0\]

OpenStudy (anonymous):

do i evaluate it as 6^(2x) - 12^x - 15 = 0 first?

OpenStudy (cruffo):

this looks like a disguised quadratic equation

OpenStudy (cruffo):

If you let \(\large u = 6^x\) then \(\large u^2 = 6^{2x}\) so \(\large u^2 - 2u - 15 = 0\)

OpenStudy (anonymous):

okay so from there... i factor it out right?

OpenStudy (cruffo):

right.

OpenStudy (anonymous):

Oh... wow. okay so for this case i keep the 2 out. and don't multiply it with 6 then?

OpenStudy (cruffo):

right - order of operations - 2 * 6^x is not equal to 12^x

OpenStudy (cruffo):

for example if x is 2 then 2*6^2 = 2*36 = 72 but 12^2 = 144

OpenStudy (anonymous):

wait so i got (u+5)(u-3) when i factored right.. then 2(6) ... huh??? :S omg i got lost again..

OpenStudy (cruffo):

lol...I was answering your side question about 2*6^x verses 12^x with an example

OpenStudy (cruffo):

you should get (u-5)(u+3) when you factor

OpenStudy (anonymous):

oh lol ok oh and ye so i then reject the negative and get 3. 6^x = 3 ln(6)^x=ln(3) xln(6)=ln(3) ... uhh... ? what now? ... wait isn't it (u+5)(u-3)? not (u-5)(u+3)

OpenStudy (cruffo):

yep, the signs...

OpenStudy (cruffo):

middle term is negative, so larger factor is negative

OpenStudy (anonymous):

wait woah. i just confused myself... -.-

OpenStudy (anonymous):

sorry! nvm what i said. ooook. soo then it would be x = ln(3)/ln(6)

OpenStudy (cruffo):

almost...but.... when you factor: \[\large u^2 - 2u - 15 = (u-5)(u+3)\] so u = 5 and u = -3 reject the negative answer and then \[\large 6^x = 5\] \[\large x = \frac{\ln5}{\ln6}\]

OpenStudy (anonymous):

iofgiohda omg ! .. i forgot to switch the signs. ok so i have another q. and it is \[4^{2x} + 9(4^x) + 14=0\] and i got two negative roots... would that mean that x = 0 and i don't do anything else to it?

OpenStudy (cruffo):

looks like no solution then

OpenStudy (anonymous):

okay thank you again. im sorry if i might be really bad at this... the notes given by the teacher are really poorly written.

OpenStudy (cruffo):

you're really not bad at this at all.

OpenStudy (anonymous):

oh well thanks haha